answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

How rare is a hole in one?

Approximate odds are 8200 to 1 on a average length par three hole tee.

What does 'accountability' mean?

When someone has "accountability", it means they are able to be depended on.

How do you to get the probability in normal distribution?

You calculate the z-scores and then use published tables.

What is the probability of getting a number greater than four when rolling a dice?

When you throw a dice there is always 6 chances of getting a number; therefore the numbers greater that 4 are: 5 and 6 (2 options)

in total there will be 2/6 chances will simplifies to 1/3

How many distinguishable permutations can be made from the letters in the word gallipolis?


In the word "gallipolis"

there are 10 letters (n =

10).

"l" appears 3 times (n1 =

3).

"i" appears 2 times (n2 =

2).

And 5 letters appear once (n3 =1, n4 =

1, n5 =1, n6 =1, n7 =1)

The number of permutations that can be made with these 10 letters is;

P =

n!/(n1!n2!n3!n4!n5!n6!n7!) =

10!/(3!∙2!∙1!∙1!∙1!∙1!∙1!) =

302 400

How do you get 1000 stickers on first in math?

type in first in math on your home page do not prees space on the top on your home page you`ll see hack prees that it will tell how much stickers you would like .

What is the fixed probability of 12?

There is insufficient information in the question to answer it. Please restate the question.

What is the probability of pulling a yellow-green marble out of the bag?

It depends on how many yellow-green marbles there are, and on how many total marbles there are. There is insufficient information in the question to answer it. Please restate the question, giving this other information.

Which shows the probability of choosing 2 white socks from a drawer having 6 white 3 black socks 3 brown socks and 8 gray socks?

I assume you are selecting two socks (one at a time) from the drawer to wear (for example).

There are 6 white + 3 black + 3 brown + 8 gray = 20 socks in all

So the probability that the first sock chosen to be white is 6/20 since there are 6 socks and 20 socks in total. 6/20 reduces down to 3/10.

The probability that the second sock chosen is also white is 5/19 since there are now only 5 white socks left to be chosen and 19 socks in total (since one sock has been taken out).

Thus the probability of both socks being white is:

probability = 3/10 x 5/19

= 3/38

What is the probability that an event will happen?

It is a measure of the likelihood of that event occurring, as a proportion of all possible outcomes.

What is the chance of getting a five in a pack of playing cards?

There are 52 cards in a deck, and 4 suits of 13 cards each.

So the probability is 1/13 = 0.077

What three coins are most probable to select out of pennies dimes and nickels?

More information is needed. You have to know how many of each coin are in the sample set in order to determine a probability of selecting any particular coin or combination of coins.

Please post a new, separate question with that information.